# Meta-UAD: A Meta-Learning Scheme for User-level Network Traffic Anomaly Detection

### v1

@@ Abstract
Accuracy anomaly detection in user-level network traffic is crucial for network security[1]. Attackers might exploit an application containing a vulnerability, jeopardizing the confidentiality, integrity, and availability of the user’s crucial information.
User-level network traffic contains sizeable new anomaly classes with few labeled samples. Those new anomaly classes possess three unique characteristics[2]. Imbalanced: compared to the existing anomaly classes, it is expensive and arduous to collect a massive amount of data onto the new anomaly class. Therefore, the sample sizes of different anomaly classes in the sampled database are often highly imbalanced. Self-similar: the new anomaly classes evolve from the existing ones and have characteristics closer to normal traffic. As our detection models adapt, so does anomaly traffic. According to the McAfee labs threats reports [^1] https://media.mcafeeassets.com/content/dam/npcld/ecommerce/en-us/docs/reports/rp-mobile-threat-report-feb-2023.pdf., most new anomaly classes are branches of existing anomaly families. Data-hungry: now more attackers are focusing on fewer but more precise targets instead of widespread invasions, so each new anomaly class has small-scale samples. According to the CrowdStrike report [^2] https://www.crowdstrike.com/global-threat-report/., near 62% attackers will remain silent until a precise target is discovered, and they use legitimate credentials and built-in tools to attack, which can avoid detection by traditional detection models.
@@
Existing anomaly detection models can be grouped into three categories (see in Fig.1). Machine-learning (ML) based models[3, 4, 5] classify data using a feature matrix where rows and columns correspond to each packet and feature values. They can combine the rules that differentiate anomaly traffic from normal traffic and achieve high-precision classification. Metric-based models[6, 7] can learn a non-linear metric to generate new anomaly samples by distance to a class prototype or a class sub-space. Simultaneously, the generative-based models[8, 9] can learn a sample generator only from seen class samples to synthesize new samples using Generative Adversarial Networks (GAN).
@@
Existing models only detect specific anomaly classes with large labeled training samples, which are hard to detect anomaly samples in user-level network traffic. 1) They heavily rely on large-scale training samples and anomaly class distribution in training datasets[10]. They only can achieve good results for specific anomaly classes (same as the training dataset). For small-scale training datasets (e.g., endpoint traffic), the rules learned by these models may be entirely inapplicable to other datasets (e.g., cloud traffic). 2) For extremely imbalanced training datasets, models can produce an inductive learning bias towards the majority classes due to overfitting[11], resulting in poor minority class detection performance. The sample generator can alleviate the data imbalance problem. Nevertheless, the sample generator uses only seen class data and does not explicitly learn to generate unseen class samples. If we detect many unseen classes, the number of synthesizing samples would be unfeasibly estimated, and this operation will drastically increase computational complexity and energy consumption. In the testing phase, when the new anomaly class is significantly different from anomaly classes in the training set, the performance of those models is still poor[12]. 3) Since the anomaly classes are constantly updated and training data cannot cover all anomaly classes, the effectiveness of existing models will decline sharply for new anomaly classes.
@@
We argue that the recent breakthroughs in meta-learning have greatly facilitated small-scale sample processing capacities. Among them, few-shot learning[13, 14, 15, 16, 17] enables the model to effectively learn information from small samples by combining limited supervised information (few labeled samples) with prior knowledge (meta-network). Meta-SGD[18] is a dominant optimization-based few-shot learning model. Motivation on this opportunity, in this paper, we propose Meta-UAD, a Meta-learning scheme for User-level network traffic Anomaly Detection, which is based on Meta-SGD. Meta-UAD uses the CICFlowMeter[19] to extract 81 flow-level statistical features and remove some invalid ones using cumulative importance ranking. Meta-UAD adopts a meta-learning training structure and learns from the collection of K-way-M-shot classification tasks, which can use a pre-trained model to adapt any new class with few samples by few iteration steps (see in Fig.1). To evaluate our scheme’s efficiency, we conduct two comparative experiments in two public network traffic datasets, and the results further demonstrate the superiority of Meta-UAD with 15% - 43% gains in the F1-score.
I Anomaly Detection with Meta-learning
I-A Problem Definition
@@ 33 Flow flag count: ACK
Backbone architecture. Our scheme Meta-UAD is general. In theory, we can use any anomaly detection network as the backbone architecture. Because the data format of network traffic is irregular, the features of network traffic have been extracted preliminarily through the data preprocessing module. For the anomaly detection module, we only need a simple backbone architecture to learn good accuracy by using our scheme. So we come up with the Deep Neural Networks (DNN[20]) as backbone architecture.
@@
II Experiments
II-A Experimental Setup
@@
Considering the data imbalanced condition in a multiclass classification problem, to efficiently judge our scheme’s performance in detecting new anomaly classes with few labeled samples, we use F1-score as the criterion (same as prior work [23, 24, 25]). We use five ML-based models as comparison algorithms: SVM [3], DT [4], RF [5], CNN [26], LSTM [27]. Similarly, we use two metric-based models and two generative-based models as comparison algorithms: Simpleshot[7], TADAM[6], Delta-encoder[8], and Mfc-gan[9].
@@
II-B Meta-UAD on M-shot anomaly detection

### v2

@@ Abstract
I Introduction
Accuracy anomaly detection in user-level traffic is crucial for network security[1]. Attackers might exploit an application containing a vulnerability, jeopardizing the confidentiality, integrity, and availability of the user’s crucial information.
User-level network traffic contains sizeable new anomaly classes with few labeled samples. Those new anomaly classes possess three unique characteristics. Imbalanced[2, 3]: compared to the existing anomaly classes, it is expensive and arduous to collect a massive amount of data onto the new anomaly class. Therefore, the sample sizes of different anomaly classes in the sampled database are often highly imbalanced. Self-similar[4, 5]: the new anomaly classes evolve from the existing ones and have characteristics closer to normal traffic. As our detection models adapt, so does anomaly traffic. According to the McAfee labs threats reports [^1] https://media.mcafeeassets.com/content/dam/npcld/ecommerce/en-us/docs/reports/rp-mobile-threat-report-feb-2023.pdf., most new anomaly classes are branches of existing anomaly families. Data-hungry[6, 7]: now more attackers are focusing on fewer but more precise targets instead of widespread invasions, so each new anomaly class has small-scale samples. According to the CrowdStrike report [^2] https://www.crowdstrike.com/global-threat-report/., near 62% attackers will remain silent until a precise target is discovered, and they use legitimate credentials and built-in tools to attack, which can avoid detection by traditional detection models.
@@ I Introduction
Existing anomaly detection models can be grouped into three categories (see in Fig.1). Machine-learning (ML) based models[8, 9, 10] classify data using a feature matrix where rows and columns correspond to each packet and feature values. They can combine the rules that differentiate anomaly traffic from normal traffic and achieve high-precision classification. Metric-based models[11, 12, 13] can learn a non-linear metric to generate new anomaly samples by distance to a class prototype or a class sub-space. Simultaneously, the generative-based models[14, 15, 16] can learn a sample generator only from seen class samples to synthesize new samples using Generative Adversarial Networks (GAN).
@@
Existing models only detect specific anomaly classes with large labeled training samples, which are hard to detect anomaly samples in user-level network traffic. 1) They heavily rely on large-scale training samples and anomaly class distribution in training datasets[17]. They only can achieve good results for specific anomaly classes (same as the training dataset). For small-scale training datasets (e.g., endpoint traffic), the rules learned by these models may be entirely inapplicable to other datasets (e.g., cloud traffic). 2) For extremely imbalanced training datasets, models can produce an inductive learning bias towards the majority classes due to overfitting[18], resulting in poor minority class detection performance. The sample generator can alleviate the data imbalance problem. Nevertheless, the sample generator uses only seen class data and does not explicitly learn to generate unseen class samples. If we detect many unseen classes, the number of synthesizing samples would be unfeasibly estimated[19], and this operation will drastically increase computational complexity and energy consumption. In the testing phase, when the new anomaly class is significantly different from anomaly classes in the training set, the performance of those models is still poor[20]. 3) Since the anomaly classes are constantly updated and training data cannot cover all anomaly classes, the effectiveness of existing models will decline sharply for new anomaly classes[21].
@@
We argue that the recent breakthroughs in meta-learning have greatly facilitated small-scale sample processing capacities. Among them, few-shot learning[22] enables the model to effectively learn information from small samples by combining limited supervised information (few labeled samples) with prior knowledge (meta-network). Meta-SGD[23] is a dominant optimization-based few-shot learning model. Motivation on this opportunity, in this paper, we propose Meta-UAD, a Meta-learning scheme for User-level network traffic Anomaly Detection, which is based on Meta-SGD. Meta-UAD uses the CICFlowMeter[24] to extract 81 flow-level statistical features and remove some invalid ones using cumulative importance ranking. Meta-UAD adopts a meta-learning training structure and learns from the collection of K-way-M-shot classification tasks (see in Fig.1). To evaluate our scheme’s efficiency, we conduct two comparative experiments in two public network traffic datasets, and the results further demonstrate the superiority of Meta-UAD with 15% - 43% gains in the F1-score.
II Anomaly Detection with Meta-learning
II-A Problem Definition
@@ 33 Flow flag count: ACK
Backbone architecture. Our scheme Meta-UAD is general. In theory, we can use any anomaly detection network as the backbone architecture. Because the data format of network traffic is irregular, the features of network traffic have been extracted preliminarily through the data preprocessing module. For the anomaly detection module, we only need a simple backbone architecture to learn good accuracy by using our scheme. So we come up with the Deep Neural Networks (DNN) as backbone architecture.
@@
III Experiments
III-A Experimental Setup
@@
Considering the data imbalanced condition in a multiclass classification problem, to efficiently judge our scheme’s performance in detecting new anomaly classes with few labeled samples, we use F1-score as the criterion. We use five ML-based models as comparison algorithms: SVM [28], DT [29], RF [25], CNN [30], LSTM [31]. Similarly, we use two metric-based models and two generative-based models as comparison algorithms: Simpleshot[12], TADAM[11], Delta-encoder[14], and Mfc-gan[15].
@@
III-B Meta-UAD on M-shot anomaly detection
@@ IV CONCLUSION
V Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants (62171057, 62471055, U23B2001), China Postdoctoral Science Foundation under Grant No. 2024M751688, Postdoctoral Fellowship Program of CPSF under Grant No. GZC20240827, and Beijing Key Lab of Networked Multimedia.